# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v1

@@ V-F Robustness of MTKD
In the previous subsection, we have demonstrated that the MTKD framework based on three teacher models can significantly improve the performance of various CD models. Therefore, we further explore the performance of MTKD with two teacher models. We set a threshold of $th=0.10$ to divide the dataset into small and large partitions and train the models $\mathcal{M}_{T_{S}}$ and $\mathcal{M}_{T_{L}}$ on these partitions, respectively. Subsequently, the student model $\mathcal{M}_{S}$ is trained using the MTKD framework. Except for the reduced number of teacher models, all other experimental settings remain unchanged. Using the original models as a baseline, Table V compares the metrics of models trained with the O-P strategy and the MTKD framework under different numbers of teacher models. The results indicate that two teacher models can still achieve significant performance improvements over the baseline. Notably, for Changer-MiT-b0 and Changer-MiT-b1, the two-teacher model setup yields even greater improvements compared to the three-teacher setup. For CGNet, the O-P strategy with two teachers achieves higher mIoU and mFscore, although the MTKD performance is lower in this case. For TTP, the performance of O-P and MTKD under the three teacher model still remains the strongest. These experiments validate the robustness of our methods in terms of the number of teachers. They also suggest that for some models, reducing the number of teachers can lead to comparable or even superior results while consuming fewer training resources.
@@
We further validated the effectiveness of the MTKD framework on the SYSU-CD dataset. Transformer-based Changer-MiT-b1, CNN-based CGNet, and FM-based TTP models are selected. As shown in Table VI, after MTKD optimization, all the models achieve higher mIoU and mFscore values, with the most significant improvement on CGNet (mIoU increased by 0.46% and mFscore increased by 0.60%). However, the O-P strategy does not yield better results compared to the original models. The visual comparison is presented in Fig. 9, illustrating the detection results of these algorithms on four image pairs. Comparing the 4th and 5th columns, it is evident that MTKD significantly reduces the false alarm rate for Changer-MiT-b1. For example, in the 1st row, large areas of false detection are eliminated, and in the 3rd row, a small false-positive patch is removed. However, the 4th row exhibits an increase in false alarms. For CGNet, MTKD’s most notable contribution is the enhancement of internal compactness in detected regions. The original TTP model already produces satisfactory detection results, but MTKD further reduces false alarms. For instance, a large patch in the 1st row and a small patch in the 3rd row are effectively eliminated. However, the 1st row also shows an increase in missed detections. Overall, MTKD demonstrated substantial effectiveness on the SYSU-CD dataset, further validating its robustness.
@@
TABLE VI: Experimental Results on SYSU-CD Test Set
@@
Method Strategy mIoU mAcc mPrecision mFscore
Changer (MiT-b1) - 75.49 84.25 86.38 82.38
O-P 75.32 83.88 85.86 82.19
MTKD 75.56 83.97 87.33 82.40
CGNet - 71.41 80.32 84.72 78.82
O-P 69.85 80.05 82.39 77.40
MTKD 71.87 81.74 84.05 79.42
TTP - 76.09 84.59 87.08 82.72
O-P 75.97 84.54 86.20 82.66
MTKD 76.11 83.93 87.77 82.77
@@ VI Conclusion
In this work, we introduce a new benchmark dataset, JL1-CD, which significantly complements existing CD datasets by offering sub-meter resolution, a wide range of change types, and a large dataset scale. We also propose the MTKD framework, which significantly enhances the performance of CD models on dual-temporal remote sensing images with different change areas, without increasing time and computational complexity during inference. Our approach effectively improves the generalization and robustness of the model across different cD scenarios. Future work will focus on developing a more universal knowledge distillation framework that not only further enhances model performance in diverse scenarios but also reduces model size and increases inference speed.

### v2

@@ V-F Robustness of MTKD
In the previous subsection, we have demonstrated that the MTKD framework based on three teacher models can significantly improve the performance of various CD models. However, the 3-teacher configuration requires training five versions of the model for a single CD algorithm: $\mathcal{M}_{O}$, $\mathcal{M}_{T_{L}}$, $\mathcal{M}_{T_{M}}$, $\mathcal{M}_{T_{S}}$, and the student model $\mathcal{M}_{S}$. To alleviate the resource consumption associated with training these models and testing under the O-P strategy, we further explore the performance of the MTKD framework with only two teacher models. We set a threshold of $th=0.10$ to divide the dataset into small and large partitions and train the models $\mathcal{M}_{T_{S}}$ and $\mathcal{M}_{T_{L}}$ on these partitions, respectively. Subsequently, the student model $\mathcal{M}_{S}$ is trained using the MTKD framework. Except for the reduced number of teacher models, all other experimental settings remain unchanged. Using the original models as a baseline, Table V compares the metrics of models trained with the O-P strategy and the MTKD framework under different numbers of teacher models. The results indicate that two teacher models can still achieve significant performance improvements over the baseline. Notably, for Changer-MiT-b0 and Changer-MiT-b1, the two-teacher model setup yields even greater improvements compared to the three-teacher setup. For CGNet, the O-P strategy with two teachers achieves higher mIoU and mFscore, although the MTKD performance is lower in this case. For TTP, the performance of O-P and MTKD under the three teacher model still remains the strongest. These experiments confirm that the O-P strategy and MTKD framework remain effective with fewer teachers. In some cases, the 2-teacher configuration even outperforms the 3-teacher setup, indicating the flexibility of our approach.
@@
We further evaluate the effectiveness of the MTKD framework on the SYSU-CD dataset. To ensure the representativeness and research significance of the experiments, we select one model from each of the three mainstream frameworks. The models are chosen with careful consideration:
  1. 1.
    TTP (FM-based) [34]: It represents the most recent and state-of-the-art FM-based model. Furthermore, it achieves the best performance on the JL1-CD dataset after optimization using our proposed MTKD framework.
  2. 2.
    Changer-MiT-b1 (Transformer-based) [14]: With only 13.355M parameters, this model achieves the second-highest mIoU and mFscore on the JL1-CD dataset, demonstrating its efficiency and effectiveness.
  3. 3.
    CGNet (CNN-based) [11]: Among all CNN-based models, CGNet has the largest parameter count and delivers the best performance, making it a strong representative of the CNN architecture.
We summarize the experimental results of these three algorithms on the SYSU-CD dataset, under different configurations of the number of teacher models, as well as their detection metrics for both change and no-change classes in Table VI. As shown in Table VI, regardless of the number of teacher models, the O-P strategy does not yield better results compared to the original models. The failure of the O-P strategy on the SYSU-CD dataset highlights its limitations. However, our MTKD approach still achieves notable success. After MTKD optimization, all models, whether with three or two teacher models, demonstrate higher mIoU and mFscore values. Notably, CGNet shows the most significant performance improvement in the three-teacher setup, while Changer-MiT-b1 and TTP perform better in the two-teacher setup. This contrasts with the test results on the JL1-CD dataset, suggesting that the optimal number of teacher models may vary depending on the dataset. This also indicates the need for further development of a method capable of automatically selecting the appropriate number of teachers based on the scenario. The Changer-MiT-b1 model with two teacher models shows the highest improvement in mIoU (+0.47%), while CGNet with three teacher models achieves the greatest improvement in mFscore (+0.60%).
In the last four columns of the table, we also present the detection metrics for both the change and no-change classes. For clarity, we show the absolute values obtained by the original models, while the results under the O-P strategy and the MTKD framework are represented as relative values, facilitating a more intuitive comparison of how our methods focus on the detection of change versus no-change areas. According to the MTKD setup with the optimal number of teachers, we observe higher improvements in the change class (e.g., the IoU and Fscore for the change class with CGNet in the three-teacher setup increased by 1.40% and 1.61%, respectively, while performance in the no-change class actually decreased). This demonstrates that MTKD is more sensitive to the detection of change areas, which is consistent with our findings on the JL1-CD dataset.
@@
TABLE VI: Experimental Results on SYSU-CD Test Set (Best and Second-Best Results are in Bold and Underlined, Respectively)
@@
Method Strategy No. of $\mathcal{M}_{T}$ mIoU mAcc mPrecision mFscore Class IoU Acc Precision Fscore
Changer (MiT-b1) - - 75.49 84.25 86.38 82.38 unchanged 87.75 93.93 92.86 91.92
changed 63.23 74.58 79.91 72.85
O-P 3 75.32 83.88 85.86 82.19 unchanged +0.07 +0.19 +0.07 +0.12
changed -0.40 -0.93 -1.12 -0.52
2 75.47 83.79 86.34 82.31 unchanged +0.27 +0.48 -0.02 +0.27
changed -0.32 -1.40 -0.08 -0.41
MTKD 3 75.56 83.97 87.33 82.40 unchanged +0.15 +0.44 -0.12 +0.08
changed +0.00 -1.02 +2.02 -0.05
2 75.96 84.31 87.03 82.76 unchanged +0.37 +0.20 +0.38 +0.29
changed +0.56 -0.09 +0.92 +0.45
CGNet - - 71.41 80.32 84.72 78.82 unchanged 86.08 93.96 90.96 90.83
changed 56.75 66.67 78.49 66.81
O-P 3 69.85 80.05 82.39 77.40 unchanged -1.59 -3.14 +1.45 -1.12
changed -1.54 +2.60 -6.12 -1.72
2 68.37 77.02 85.79 75.58 unchanged -0.61 +0.08 -0.66 -0.51
changed -5.48 -6.66 +2.79 -5.97
MTKD 3 71.87 81.74 84.05 79.42 unchanged -0.49 -2.18 +1.57 -0.42
changed +1.40 +5.03 -2.91 +1.61
2 71.75 80.64 84.69 79.21 unchanged +0.22 -0.01 +0.10 +0.18
changed +0.44 +0.66 -0.17 +0.59
TTP - - 76.09 84.59 87.08 82.72 unchanged 88.30 94.46 93.08 92.26
changed 63.88 74.71 81.08 73.17
O-P 3 75.97 84.54 86.20 82.66 unchanged -0.41 -0.79 +0.42 -0.26
changed +0.18 +0.71 -2.18 +0.15
2 75.83 83.71 86.69 82.83 unchanged -0.44 +0.24 -0.97 -0.38
changed -0.09 -1.99 +0.20 +0.60
MTKD 3 76.11 83.93 87.77 82.77 unchanged +0.05 +0.65 -0.73 +0.03
changed +0.00 -1.96 +2.11 +0.07
2 76.53 85.54 86.22 83.29 unchanged -0.26 -0.85 +0.41 -0.19
changed +1.14 +2.75 -2.13 +1.34
The visual comparison is presented in Fig. 9, illustrating the detection results of these algorithms on four image pairs. Comparing the 4th and 5th columns, it is evident that MTKD significantly reduces the false alarm rate for Changer-MiT-b1. For example, in the 1st row, large areas of false detection are eliminated, and in the 3rd row, a small false-positive patch is removed. However, the 4th row exhibits an increase in false alarms. For CGNet, MTKD’s most notable contribution is the enhancement of internal compactness in detected regions. The original TTP model already produces satisfactory detection results, but MTKD further reduces false alarms. For instance, a large patch in the 1st row and a small patch in the 3rd row are effectively eliminated. However, the 1st row also shows an increase in missed detections. Overall, MTKD demonstrated substantial effectiveness on the SYSU-CD dataset, further validating its robustness.
@@ VI Conclusion
In this work, we introduce a new benchmark dataset, JL1-CD, which significantly complements existing CD datasets by offering sub-meter resolution, a wide range of change types, and a large dataset scale. We also propose the MTKD framework, which significantly enhances the performance of CD models on dual-temporal remote sensing images with different change areas, without increasing time and computational complexity during inference. Our approach effectively improves the generalization and robustness of the model across different CD scenarios. Future work will focus on the adaptive selection of the number of teacher models and partition thresholds based on different scenarios, in order to enhance the flexibility and performance of the MTKD framework. Additionally, the evaluation of the MTKD framework on datasets with more complex and extreme change conditions, such as those affected by seasonal variations, severe weather events, or rapid urbanization, will be a critical direction for future research. This will help assess the robustness and adaptability of the framework under challenging real-world conditions.